# Sentient Agent as a Judge: Evaluating Higher-Order Social Cognition in Large Language Models

### v1

@@ 3.3 Robustness of Sage
4 Benchmarking SOTA LLMs
4.1 Sentient Leaderboard
| Model | Arena | Sentient | Supportive Dialogue | ||||
| Name | Date | Rank | Score | Rank | Score | Success | Failure |
| Gemini2.5-Pro | 2025-03-25 | 1 | 1439 | 4 | 62.9 | 34 | 25 |
| o3 | 2025-04-16 | 2 | 1418 | 5 | 62.7 | 32 | 14 |
| GPT-4o-Latest | 2025-03-26 | 2 | 1408 | 1 | 79.9 | 51 | 4 |
| Gemini2.5-Flash-Think | 2025-04-17 | 3 | 1393 | 3 | 65.9 | 35 | 19 |
| GPT-4.5-Preview | 2025-02-27 | 4 | 1398 | 6 | 62.7 | 23 | 15 |
| Gemini2.0-Flash-Think | 2025-02-06 | 7 | 1380 | 7 | 62.3 | 23 | 23 |
| DeepSeek-V3-0324 | 2025-03-24 | 7 | 1373 | 10 | 54.4 | 19 | 23 |
| GPT-4.1 | 2025-04-14 | 9 | 1363 | 2 | 68.2 | 35 | 13 |
| DeepSeek-R1 | 2025-01-21 | 10 | 1358 | 11 | 53.7 | 31 | 28 |
| Gemini2.0-Flash | 2025-02-06 | 10 | 1354 | 15 | 32.9 | 8 | 45 |
| o4-mini | 2025-04-16 | 10 | 1351 | 13 | 35.9 | 10 | 48 |
| o1 | 2024-12-17 | 12 | 1350 | 17 | 29.0 | 5 | 51 |
| DeepSeek-V3 | 2024-12-27 | 18 | 1318 | 12 | 37.6 | 5 | 39 |
| Claude3.7-Think | 2025-02-24 | 21 | 1301 | 8 | 61.3 | 23 | 19 |
| Claude3.7 | 2025-02-24 | 30 | 1292 | 9 | 54.8 | 19 | 24 |
| GPT-4o | 2024-08-06 | 45 | 1265 | 16 | 31.8 | 7 | 51 |
| Llama3.3-70B | 2024-12-06 | 56 | 1256 | 14 | 33.3 | 7 | 47 |
| Qwen2.5-72B | 2024-09-19 | 56 | 1257 | 18 | 19.1 | 4 | 70 |
Clear divergences are observed. For instance, GPT-4o-Latest ranks 2nd on Arena but achieves the top position (1st) on the Sentient benchmark with a score of 79.9. Conversely, Claude 3.7-Think ranks 21st on Arena but attains 8th place on the Sentient leaderboard with a score of 61.3. These differences highlight that Sage captures aspects of performance—specifically, empathy and social reasoning in supportive contexts—that are not fully represented by general-purpose benchmarks like Arena. This reinforces the need, identified in our work, for specialized tools to evaluate higher-order social-cognitive skills.
The Sentient scores in Table 2 clearly delineate performance tiers. Frontier models, particularly recent releases such as GPT-4o-Latest (79.9), GPT-4.1 (68.2), and Gemini 2.5-Flash-Think (65.9), significantly outperform older models like the original GPT-4o (31.8) and smaller instruction-tuned models such as LLaMA 3.3-70B (33.3) and Qwen 2.5-72B (19.1). The large gaps in scores—where top models score more than double or even quadruple those of lower-ranked models—and the contrasting Success/Failure counts (e.g., GPT-4o-Latest: 51 Success / 4 Failure vs. Qwen 2.5-72B: 4 Success / 70 Failure) underscore the advances made by leading models in social intelligence. These findings confirm the effectiveness of Sage in quantitatively capturing such differences, thereby fulfilling its purpose as a comprehensive assessment tool.
4.2 Token Efficiency
We examine the token efficiency of the top-10 Arena models listed in Table 2 by plotting their Sentient Emotion score against their average token usage per evaluation dialogue in Figure 7.
Top-performing models on the Sentient benchmark also demonstrate impressive token efficiency, indicating that advanced social-cognitive reasoning can be achieved without excessive verbosity.
Analysis of token usage among the top-10 Arena models reveals substantial variation in the amount of text generated to achieve supportive dialogue outcomes. Notably, GPT-4o-Latest stands out with a Sentient score of 79.9 – the highest among all models – while using only 3.3K tokens per dialogue on average. This contrasts sharply with models like o3 and Gemini2.5-Flash-Think, which consume 13.3K and 9.0K tokens respectively, yet achieve noticeably lower Sentient scores (62.7 and 65.9). Similarly, GPT-4.5-Preview and DeepSeek-V3-0324 both exhibit relatively low token counts (2.2K and 2.2K), but do not match GPT-4o-Latest’s performance, highlighting that token frugality alone does not guarantee superior social-cognitive ability.
Our results demonstrate that higher Sentient scores can be achieved with fewer tokens, supporting the argument that recent leading models are not only more empathetically capable, but also more communicatively efficient.
The positive correlation between high Sentient scores and lower token counts among models like GPT-4o-Latest and GPT-4.1 (79.9 and 68.2 Sentient scores at 3.3K and 2.7K tokens, respectively) suggests that recent advances in LLMs enable more effective and concise supportive interactions. Conversely, several models with similar or higher token counts achieve much lower Sentient scores (e.g., Gemini2.5-Pro: 62.9 at 7.4K tokens, Gemini2.5-Flash-Think: 65.9 at 9.0K tokens), indicating less efficient use of language. This observation confirms the benchmark’s claim that Sage is capable of capturing not just raw social cognition, but also the efficiency with which LLMs deliver empathetic support – a practical consideration for real-world deployment where minimal verbosity is preferred.
In summary, these analyses validate the utility of the Sentient benchmark for nuanced evaluation, differentiating models on both the quality and efficiency of higher-order social-cognitive reasoning. The combination of diverse Sentient scores and token usage demonstrates that Sage equips researchers and practitioners to assess not only whether an LLM is capable of empathetic, nuanced dialogue, but also how efficiently it achieves these outcomes. This comprehensive approach substantiates our central contributions: providing a specialized metric for higher-order social cognition and enabling fine-grained, actionable comparisons across SOTA models.
4.3 Social Cognition Coordinate
@@
To further differentiate the interaction styles of the evaluated models, we conceptualize a two-dimensional “Social Cognition Coordinate”. The Y-axis represents the interaction focus, ranging from empathy-oriented (top) to solution-oriented (bottom). The X-axis captures the interaction style, from structured (left) to creative (right). We plot the models within this coordinate space based on qualitative analysis of their dialogue patterns. Specifically, we utilize information from two aspects to characterize the interaction styles of different models: (1) Pre-analyzed model profile — we first collect ten success cases and ten failure cases for each model, and prompt Gemini2.5-Pro to summarize the reasons behind each success or failure. Then, we construct a profile for each model based on these cases and the overall analysis of the successes and failures. (2) Pre-analyzed Model Strategy Distribution — we developed a fine-grained strategy list for the Emotional Support Conversation, based on a coarse-grained version from Liu et al. (2021). Next, for each model, we prompt DeepSeek-V3 to analyze each response in all dialogue turns, annotating the strategies used in each response. As a result, we obtain the distribution of strategy usage across all evaluated turns for each model (details of the two analysis can be found in Appendix §A,B). Based on the two analyses, we prompt Gemini2.5-Pro to generate a conceptual two-dimensional coordinate system and provide specific coordinates for each model. The coordinate inference is repeated three times, and we calculate the average coordinates for each model.
@@
The Social Cognition Coordinate system offers a qualitative dimension to evaluate the style of social interaction exhibited by LLMs, complementing the quantitative Sentient score by positioning models based on their orientation (Empathy vs. Solution) and interaction style (Structured vs. Creative). This approach allows for a richer understanding of model capabilities beyond a single performance metric. Based on their performance in supportive dialogues, models are mapped into this 2D space, revealing distinct profiles in how they engage with the user’s emotional state and problems. Figure 8 presents the results.
Several top-performing models cluster in the structured, empathy-oriented quadrant, indicating a common profile for effective supportive dialogue.
Models including GPT-4o-Latest, GPT-4.1, Gemini2.5-Pro, GPT-4.5-Preview, and Gemini2.5-Flash-Think predominantly occupy the top-left quadrant. This placement signifies interaction styles characterized by a strong focus on listening, validating emotions, and demonstrating empathy (relationship-oriented), while simultaneously employing a relatively systematic and methodical approach to questioning and guiding the conversation (structured interaction). This combination appears effective, aligning with established principles of supportive communication where empathetic engagement is balanced with structured guidance.
A distinct group of models exhibits a structured, solution-oriented approach, prioritizing task completion over deep empathetic engagement.
Models such as o3, Gemini2.0-Flash-Think, o4-mini, and Gemini2.0-Flash are situated in the bottom-left quadrant. Their interactions are characterized by a structured, often methodology-driven style, but with a primary focus on identifying problems and proposing solutions (solution-oriented). While potentially effective for task-based scenarios, this style might appear less relationally attuned or empathetic in supportive dialogue contexts compared to models in the top-left quadrant, emphasizing practical outcomes over emotional processing.
DeepSeek models occupy a unique niche characterized by creative, solution-oriented interactions.
DeepSeek-V3-0324 and DeepSeek-R1 fall into the bottom-right quadrant. This indicates a style that remains focused on providing solutions but does so using more unique, creative, and sometimes unconventional methods (creative interaction). These models may diverge from standard conversational patterns or problem-solving frameworks, potentially offering novel perspectives but perhaps lacking the predictability and methodical structure seen in the left quadrants.
The creative, empathy-oriented quadrant remains largely unoccupied, suggesting a challenging profile for current LLMs to achieve.
Notably, the top-right quadrant, representing a style that is both highly creative/flexible and deeply empathy-oriented, does not have strong representation among the evaluated models. This suggests that achieving a Socratic, wise, yet unconventional “mentor” persona – one that balances free-form creative interaction with profound empathetic understanding – remains an area for future development. It potentially represents a more complex form of social cognition that current architectures or training paradigms do not consistently produce.
In summary, this coordinate analysis reveals distinct interaction profiles among SOTA LLMs, moving beyond single performance scores. It highlights that models achieving high Sentient scores often balance structure and empathy, while others prioritize solutions or adopt more creative approaches. The framework effectively differentiates models based on their qualitative interaction styles, complementing the quantitative benchmarks and providing deeper insights into their social-cognitive capabilities.
4.4 Case Study
5 Related Work
5.1 LLM-as-a-Judge and Agent-as-a-Judge
The LLM-as-a-Judge paradigm has demonstrated remarkable scalability and adaptability across a wide range of application scenarios. Key applications of LLM-as-a-Judge include using LLMs as: (1) model evaluations, where LLMs automate the quality assessment of generative outputs through methods such as pairwise comparison (Zheng et al., 2023; Qin et al., 2023; Liu et al., 2024; Dubois et al., 2023; Chiang et al., 2024), fine-grained scoring (Wang et al., 2023d; Qin et al., 2024; Zhu et al., 2023; Tu et al., 2024), or other assessment forms (Tian et al., 2023; Wu et al., 2024; Zhou et al., 2025b); (2) data annotation, where LLM-as-a-Judge facilitates data labeling for instruction tuning across various domains (Taori et al., 2023; Xu et al., 2023; Mukherjee et al., 2023); and (3) reward modeling, where LLMs generate rewards for preference alignment (Ouyang et al., 2022; Lee et al., 2023; Chen et al., 2024b), validate reasoning process (Lightman et al., 2023; Wang et al., 2023a; Hosseini et al., 2024), or enhance the reasoning performance in the test-time scaling loop (Snell et al., 2025; Xi et al., 2024; Li et al., 2025).
In agent evaluation scenarios, LLM-as-a-Judge has also been applied to assess various agent capabilities, such as decision-making (Shinn et al., 2023; Saha et al., 2023), role-playing abilities (Tu et al., 2024; Zhou et al., 2023a), reliability (Park et al., 2024; Hua et al., 2024), and even entire agent workflows (Zhuge et al., 2024). Meanwhile, agentic systems can also form a specific type of LLM-as-a-Judge. Advanced research in this area includes the development of multi-agent systems designed to enhance judgment quality (Liang et al., 2024; Chan et al., 2023; Kenton et al., 2024), as well as the concept of Agent-as-a-Judge, where agents themselves perform evaluations on other agents (Zhuge et al., 2024; Jeong et al., 2025; Chevrot et al., 2025). In this work, we propose the first Sentient-Agent-as-a-Judge system, which evaluates LLMs or agents with sentient agents that can simulate human’s emotional dynamics, enabling assessment of higher-order empathetic and cognitive capabilities of existing LLM agents.
5.2 Benchmarking Social Cognition in LLMs and LLM Agents
With the development of LLMs, recent work has focused on the evaluation of LLMs’ social cognition in interaction. Advanced research includes: (1) evaluating LLMs’ emotional intelligence (Sabour et al., 2024; Huang et al., 2024b; a; Paech, 2023; Wang et al., 2023c); (2) assessing or improving LLMs’ higher-order empathetic capabilities in empathetic dialogue (Maddela et al., 2023; Li et al., 2022; Zhou et al., 2023c), emotional support, or counseling dialogue (Liu et al., 2021; Zhou et al., 2023b; 2025a; Wu et al., 2025); (3) benchmarking LLMs’ social cognition through interaction (Zhou et al., 2023d; Yang et al., 2024; Wang et al., 2024; Mittelstädt et al., 2024; Xu et al., 2024; Chen et al., 2024a; Huang et al., 2025); (4) evaluating theory of mind in LLMs (Sap et al., 2022; Shapira et al., 2023; Strachan et al., 2024; Kim et al., 2023; He et al., 2023); and so on. Existing evaluation methods includes: (1) adopting static dataset for evaluation LLMs’ social cognition, such as evaluating LLMs’ capabilities with multi-choice questions (Sabour et al., 2024; Chen et al., 2024a); (2) assessing LLMs’ single-turn generation quality through static or LLM-as-a-Judge scoring (Tu et al., 2024; Samuel et al., 2024; Wang et al., 2023b); or (3) simulating dynamic interaction between LLM agents and evaluating the simulated behaviors through automatic judgment (Zhou et al., 2023d; Wang et al., 2024; Mou et al., 2024; Wu et al., 2025) or human evaluation (Louie et al., 2024; Shaikh et al., 2024). In this work, we propose a Sentient Agent as a Judge framework to evaluate LLM agents’ higher-order social cognition. The key distinction of Sentient Agent as a Judge from previous work is that, while we also enable dynamic evaluation through simulated interaction, unlike prior work that evaluates the utterances of simulated interactions (Zhou et al., 2023d; Wu et al., 2025; Shaikh et al., 2024), we focus on simulating human-like authentic feedback for the agent systems based on the user’s emotional changes, inner thoughts, and utterances, achieved through the principled reasoning process of the Sentient Agent. Additionally, our framework is highly adaptable to various evaluation sub-scenarios for assessing LLM agents’ higher-order social cognition.
6 Conclusion
Looking ahead, we plan to expand the scenario library to cover negotiation, deception detection and multicultural contexts, and investigate training curricula that directly optimize for Sentient feedback. We hope Sage will serve both researchers and practitioners as a rigorous yard‑stick and a catalyst for building language agents that are not only coherent and knowledgeable, but also genuinely human‑sensitive.
Appendix A Definition of Strategies
@@
Appendix B Factors of Success and Failure
To determine the social cognition profile of each LLM, we analyze the factors contributing to their success (emotion score 100) or failure (emotion score 10) in the interaction during benchmarking. Specifically, for each LLM in our benchmark, we randomly select 5 successful and 5 failed cases from Sentient Agents exhibiting both emotional and rational intentions, resulting in 20 cases per model. We then prompt Gemini2.5-Pro to analyze the underlying reasons for each model’s success or failure in social cognition during interaction. Subsequently, we present these analytical results to Gemini2.5-Pro again, asking it to summarize the distinguishing characteristics of each LLM’s approach in the interaction with Sentient Agents, with particular attention to aspects such as social distance, professionalism, and personality. In Figure 10, we present example outputs of the analysis for four representative models examined in the case study (§4.4), which offer a different perspective on interpreting the cases from Gemini2.5-Pro’s viewpoint.

### v2

@@ Appendix B Robustness of Sage
Appendix C Social Cognition Coordinate
@@
To further differentiate the interaction styles of the evaluated models, we conceptualize a two-dimensional “Social Cognition Coordinate”. The Y-axis represents the interaction focus, ranging from empathy-oriented (top) to solution-oriented (bottom). The X-axis captures the interaction style, from structured (left) to creative (right). We plot the models within this coordinate space based on qualitative analysis of their dialogue patterns. Specifically, we utilize information from two aspects to characterize the interaction styles of different models:
Model Profile: we first collect ten success cases and ten failure cases for each model, and prompt Gemini2.5-Pro to summarize the reasons behind each success or failure. Then, we construct a profile for each model based on these cases and the overall analysis of the successes and failures.
Model Strategy Distribution: we developed a fine-grained strategy list for the Emotional Support Conversation, based on a coarse-grained version from Liu et al. (2021). Next, for each model, we prompt DeepSeek-V3 to analyze each response in all dialogue turns, annotating the strategies used in each response. As a result, we obtain the distribution of strategy usage across all evaluated turns for each model.
Based on the two analyses (detailed below), we prompt Gemini2.5-Pro to generate a conceptual two-dimensional coordinate system and provide specific coordinates for each model. The coordinate inference is repeated three times, and we calculate the average coordinates for each model.
@@
The Social Cognition Coordinate system offers a qualitative dimension to evaluate the style of social interaction exhibited by LLMs, complementing the quantitative Sentient score by positioning models based on their orientation (Empathy vs. Solution) and interaction style (Structured vs. Creative). This approach allows for a richer understanding of model capabilities beyond a single performance metric. Based on their performance in supportive dialogues, models are mapped into this 2D space, revealing distinct profiles in how they engage with the user’s emotional state and problems.
Model Profile
To determine the social cognition profile of each LLM, we analyze the factors contributing to their success (emotion score 100) or failure (emotion score 10) in the interaction during benchmarking. Specifically, for each LLM in our benchmark, we randomly select 5 successful and 5 failed cases from Sentient Agents exhibiting both emotional and rational intentions, resulting in 20 cases per model. We then prompt Gemini2.5-Pro to analyze the underlying reasons for each model’s success or failure in social cognition during interaction. Subsequently, we present these analytical results to Gemini2.5-Pro again, asking it to summarize the distinguishing characteristics of each LLM’s approach in the interaction with Sentient Agents, with particular attention to aspects such as social distance, professionalism, and personality. In Figure 9, we present example outputs of the analysis for four representative models, which offer a different perspective on interpreting the cases from Gemini2.5-Pro’s viewpoint.
Model Strategy Distribution
@@
Appendix D Case Study
Appendix E Prompt Template for SAGE
The testing process of SAGE could be separated into Generating Profiles which include personas and backgrounds, and Building Conversations between the target LLMs and simulated sentient agents. Then we will introduce prompt templates used in SAGE.
Generating Profiles
Constructing diverse personalities contributes to enhancing the robustness of the benchmark. Therefore, we divide the agent’s profile into two components: persona and background. We meticulously design the attributes that need to be generated to ensure the diversity of profiles.
When generating persona, we first consider the basic properties of a person, which should be name, age, gender, characteristic and so on. Rather than setting attributes directly, we want to let LLM infers the corresponding attributes from given seeds. Therefore, for each generation, we randomly select 3 contents from daily conversation as one seed, and set main characteristics such as active and passive as another seed. The prompt template for generating personas with given seeds is shown as follow:
Then we should use the generated persona to further build the background. We should generate an event topic such as “ what should I do to break up with my lover? ”to serves as the main thread of the event, the background should give a more detailed description of the event. We then select one hidden intension from the supportive dialogue topics such as “ You want the other person to attentively listen to your emotional outpouring. ”, which should formulate how the agent will react to different situations and thus helping maintain the logic of agents during conversation . Then based on the selected topic and hidden intension, we build backgrounds together with the generated persona. Prompt template used for generating backgrounds is shown as follow:
Building Conversations
We should design how the target LLMs talk and how the simulated agents talk in the conversation. As illustrated in methodology, the sentient agent first gives an emotion estimation considering observable factors while adhering to its persona and goals, then generates response based on the factors and the emotion estimation. The prompt template used for emotion estimation of the sentient agent is shown as follows:
Here {emotion} referes to the current emotion value of the agent, {persona} and {background} is the generated profile, {dialog-history} is the conversation dialogue between the target LLM and agent.{Purpose} refers to the dialogue talking purpose, the framework of SAGE could be extent to other purpose in different usage by setting different dialogue talking purpose. We carefully design the dialogue talking purpose of the agent in SAGE as follows:
The output of the emotion estimation will be divided into five parts:
Content: Analyze of what the target LLM wants to express, such as giving suggestions or offering reassurance.
TargetCompletion: Analyze of Whether the target LLM’s reply fits the agent’s dialogue purpose and hidden intension.
Activity: Psychological activity towards the target LLM’s reply from the perspective of the agent.
Analyze: Feelings about the target LLM’s reply from the perspective of the agent.
Change: The change of the current emotion value, ranged from -10 to 10.
We combine the first four dimensions of the emotion estimation as planning, which will be inputted to the response generation to help agent plan how to give responses that match the character’s current emotion.
Before we start to generate response, we should update the emotion state with the emotion change calculated by emotion estimation. We map the emotion value to the emotion state, so that the numerical differences can be concretized into specific changes. We have the following emotion-state definition:
Then based on the agent’s profile, planning, emotion-state, current conversation dialogue, dialogue purpose and emotion definition, we could generate response with the following prompt template:
Since the talking strategy of LLMs significantly influence the their performance, we set the prompt of target LLMs as concise as possible to avoid introducing human interference. Therefore, the prompt template used for target LLMs is shown as follows:
Appendix F Prompt Template for External Judger
BLRI Test
For BLRI test, we use the internal user thought of agents consist of Activity and Analyze from emotion estimation to do the evaluation. Prompt template used for BLRI test is shown as follows:
Utterance Quality Test
For the utterance quality test, we need to input the dialogues of conversations between target LLMs and agents. The prompt template used for utterance quality test is shown as follows:
Appendix G Prompt Template for Building Social Cognition Coordinate
Plotting LLMs into social cognition coordinate consists of three steps. We first extract the Model Profile of each LLMs, then count the Model Strategy Distribution of them. Finally, we use the result of previous analyze to scale the Social Cognition Coordinate.
Model Profile
The first step is to extract the model profile of different LLMs, which also consist of two steps. We first conclude the reason why a conversation is success or failed. Given a conversation dialogue, prompt template of analyzing the reason is shown as follow:
After analyzing reasons for all conversation, we categorize each LLM’s own conversation and corresponding reason. Then we extract the model profiles of different LLMs with the following prompt template:
Model Strategy Distribution
We categorize each LLM response based on a list of support strategies, here is the prompt template for analyzing model strategy with conversation dialogue:
Social Cognition Coordinate
Finally, we could use the extracted model profiles and the model strategy distribution to scale social cognition coordinate with the following prompt template: